The breaking strengths of cable
1
Expert's answer
2020-03-03T17:10:37-0500

The breaking strengths of cables produced by manufacturer have a mean of 1800 Kg and a standard deviation of 100 Kg. By a new technique in the manufacturing process, it is claimed that breaking strength can be increased. To test this claim, a sample of 50 cables is tested and it is found that the mean breaking strength is 1850Kg. Can we support the claim at the 0.01 significance level? 

Consider Z-test for one population mean "(\mu)," with known population standard deviation "(\sigma)."

The provided sample mean is "\bar{X}=1850" and the known population standard deviation is "\sigma=100," and the sample size is "n=50."

The following null and alternative hypotheses need to be tested:

"H_0: \mu=1800"

"H_1: \mu>1800"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.01," and the critical value for a right-tailed test is "z_c=2.33."

The rejection region for this right-tailed test is "R=\{z:z>2.33\}"

The z-statistic is computed as follows:


"z={\bar{X}-\mu \over \sigma/\sqrt{n}}={1850-1800 \over 100/\sqrt{50}}\approx3.535534"

Since it is observed that "z=3.535534>z_c=2.33,"  it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value is "0.000204," and since "p=0.0002<0.01," it is concluded that the null hypothesis is rejected.

It is concluded that the null hypothesis "H_0" is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is greater than 1800, at the 0.01 significance level.

We conclude that a new technique can improve the break strength.